Show that (A -B) U B = A is not necessarily true; that is, find sets A and B for which it is false.
1
Expert's answer
2021-10-11T06:59:56-0400

"A-B=\{x:x\epsilon A,x\cancel{\epsilon}B\}"

Let "A=\{1,2,3,4\}"

"B=\{3,4,5,6\}"

"A-B=\{1,2\}"

"(A-B)UB=\{1,2\}U\{3,4,5,6\}"

"=\{1,2,3,4,5,6\}"

"\not =\{1,2,3,4\}=A"

"\therefore (A-B)UB\not =A"

So the given statement is false.